### v1

# ACC Saturator: Automatic Kernel Optimization for Directive-Based GPU Code

Abstract.
Automatic code optimization is a complex process that typically involves the application of multiple discrete algorithms that modify the program structure irreversibly. However, the design of these algorithms is often monolithic, and they require repetitive implementation to perform similar analyses due to the lack of cooperation. To address this issue, modern optimization techniques, such as equality saturation, allow for exhaustive term rewriting at various levels of inputs, thereby simplifying compiler design.
In this paper, we propose equality saturation to optimize sequential codes utilized in directive/based programming for GPUs. Our approach simultaneously realizes less computation, less memory access, and high memory throughput. Our fully/automated framework constructs single/assignment forms from inputs to be entirely rewritten while keeping dependencies and extracts optimal cases. Through practical benchmarks, we demonstrate a significant performance improvement on several compilers. Furthermore, we highlight the advantages of computational reordering and emphasize the significance of memory/access order for modern GPUs.
1. Introduction
Over the last decade, the increasing popularity of computational accelerators in high-performance computing (HPC) has significantly impacted the programming landscape. The emergence of complex, unique, yet highly efficient architectures has posed a challenge for scientists, who must exert additional effort to take advantage of these architectures. Among the fastest computers, Graphics Processing Units (GPUs) are the most pervasive accelerators (TOP500.org, 2022) and are inherently parallel, requiring software to expose concurrent calculations. Traditional optimization techniques may not provide performance benefits in such scenarios. Furthermore, with the approach of the end of Moore’s Law (Moore, 1998), accelerators have become increasingly domain-specific, and compilers have become essential for the automatic utilization of supercomputing resources.
Although coding with low-level languages is typically the most profitable approach, utilizing them often involves complex data dependencies. As a result, adapting vendor/specific languages like CUDA (NVIDIA Corporation, 2022) and OpenCL (The Khronos Group Inc, 2023) requires extensive knowledge of the application and creates separate code entities just for computation offloading, which increases programming costs. To address this challenge, current GPU accelerators support programming with abstract models. OpenACC (Organization, 2011) and OpenMP (ARB, 1997) have emerged as popular models that alleviate the difficulty of accelerator use by providing code directives to existing languages. These directives allow users to specify compute/intensive parts of the original code, which compilers may automatically translate into accelerator code. At runtime, offloading is realized without any user intervention.
Directive-based code typically maintains a general programming style and remains unspecialized until compilation. It is the responsibility of compilers to handle the computation, the redundancy of memory accesses, and the order of those while considering the nature of accelerators. However, code rewriting while preserving semantics may be challenging. Current compilers apply optimization for each metric sequentially, which limits opportunities for improved performance. In particular, user-specified directives fix loop structures, so little work has been done on optimising code under directives other than relying on compiler techniques used for sequential programs (Barua et al., 2018; Matsumura et al., 2021). As a result, more complex approaches are needed to improve the performance of directive-based code.
This paper proposes the use of a modern optimization technique, equality saturation (Tate et al., 2009), to fine-tune directive/based GPU code. Our tool, ACC Saturator444The source code will be available upon publication., achieves less computation, fewer memory accesses, and high throughput at the same time, while easily integrating into the compilation of both OpenACC and OpenMP compilers. ACC Saturator performs kernel optimization by passing programs through an e-graph, a graph structure for equality saturation. Unlike other optimization methods, our approach neither transforms the abstract syntax trees nor it changes directives. Despite this, we attain significant performance improvements of up to 2.23x with the NVHPC compiler and 5.08x with GCC. We present a detailed performance analysis that highlights the benefits of each optimization on the state-of-the-art GPU architecture, NVIDIA Tesla A100. Our contributions are four-fold:
We develop a fully automated OpenACC/OpenMP framework for equality saturation.
We combine static single-assignment form with e-graph to optimize directive-based code.
We demonstrate that our approach provides significant performance opportunities for GPUs.
We provide a detailed analysis of kernel performance and the effectiveness of optimization techniques using the latest NVIDIA GPU architecture for HPC.
The rest of the paper is organized as follows. Section 2 provides background knowledge on GPUs, directive-based programming, code optimization, and equality saturation. Section 3 presents a high-level overview of our work. Sections 4 and 5 explain our approach to represent general sequential code for equality saturation and optimizations. Section 6 describes our novel code generation technique from e-graphs towards high/throughput GPU execution. In Section 7, we describe our experimental methodology. Section 8 presents the results of our evaluation along with in-depth analysis on each technique. Section 9 discusses related work. Section 10 provides concluding remarks.
2. Background
This section provides an overview of GPUs, directive-based programming, and equality saturation.
2.1. GPUs
Parallel computing is an assemblage of sequential execution, with means of intercommunication and mutual effects on efficiency. Such architectures accommodate simultaneous operations on multiple different data in uniform processors that are adjacent to each other. As an instance, a Graphics Processing Unit (GPU) runs hundreds of thousands of threads at the same time on around a hundred of streaming multiprocessors (SM). One of the most recent GPUs, NVIDIA Tesla A100, features 108 SMs and each SM contains 32 FP64 cores along with multiple LD/ST units (NVIDIA Corporation, 2020).
The primary benefit of GPU execution stems from overlapping utilization of both computational and memory units. Although GPUs provide high memory bandwidth, the global memory access latency is also higher than that of CPUs (Jia et al., 2018, 2019; Velten et al., 2022). Therefore, optimal throughput may be attained by covering memory requests with computational execution and hiding the latency of data movement. Additionally, the order of memory accesses is essential due to the memory hierarchy, because the distance among accesses by neighboring threads often results in limited bandwidth (Zohouri and Matsuoka, 2019).
2.2. Directive-Based Programming
Decomposing sequential execution into parallel entities is challenging. Routine calls often create data dependencies across multiple files, hindering the extraction of compute-intensive code. Even with extracted code, loops often possess intricate access patterns on different structures that challenge automatic parallelization for efficiency. Directive-based programming models, such as OpenACC (Organization, 2011) and OpenMP (ARB, 1997), alleviate the burden of introducing additional code for hardware acceleration. Both OpenACC and OpenMP provide code directives on the de facto standard languages for scientific applications (C/C++/ Fortran) to specify offloading parts as compute kernels with the declaration of explicit parallelism, enabling seamless automatic accelerator use.
Listing 1 provides an example of OpenACC code that utilizes code directives to specify the parallelism of loops. The kernels directive guards a sequence of kernels, while the loop directive explicitly sets the parallelism of a loop. Users may also specify a single kernel region by using the parallel directive (not shown in the listing). OpenACC’s parallelism consists of three levels of abstraction: gang, worker and vector (in coarse to fine order). For instance, on the NVHPC compiler, both the top i and the middle j loops feature gang parallelism with the degree of the trip count cy and 16, respectively, and these are distributed over thread-blocks on GPUs. The middle j loop exposes additional vector parallelism and launches 256 threads to execute the inner statements. The blue comments show an equivalent form in OpenMP, which does not allow the reuse of parallelism across nested loops. Loop parallelism and data transfers may be implicit, and the compiler automatically solves data dependencies and sets optimal parameters. The NVHPC compiler generates embarrassingly parallel code for GPUs from directive code. On the other hand, GCC utilizes a principal-agent model for both OpenACC and OpenMP, and Clang follows the same approach for OpenMP (Matsumura et al., 2023).
2.3. Equality Saturation
Compiler techniques have been critical in utilizing the best resources available for program execution on the target architecture for the last half-century (Aho et al., 2006; Wolfe et al., 1995). Various features of hardware, including registers, cache systems, pipelines, and parallelization, rely on the compiler’s efforts for performance, enabling target-specific optimization under resource constraints. However, compilers tend to miss a holistic view of performance opportunities and require additional discrete efforts for each discrete optimization mechanism (Rawat et al., 2018). Equality saturation, a state/of/the/art technique for compilers (Tate et al., 2009), defines a set of rewriting rules over the accumulation of equal expressions of target code. The rewriting continues until the expression gets saturated, finding no other forms, or hits a time or size limitation, attaining optimal codes based on a cost model that considers the entire computation.
A dedicated graph structure called e-graph accepts the accumulation of equal expressions, while sharing redundant code blocks over expressions. The graph structure consists of e-classes, groups of equal e-nodes. One e-node may point to e-classes as its subexpressions, and the accumulation works by extending or merging e-classes in accordance with rewriting rules, while preserving the relationship between parental e-nodes and child e-classes (Willsey et al., 2021; de Moura and Bjørner, 2007). To extract optimal solutions, each e-class selects one contained e-node with a minimum cost, while optionally counting common expressions only once and completing common subexpression elimination (CSE).
3. Methodology Overview
ACC Saturator is the implementation of our proposal of equality saturation for directive-based code. We provide a convenient command/line tool that wraps normal C-compiler invocation and replaces the original inputs with saturated codes.
Figure 1 provides an overview of our work. ACC Saturator optimizes the sequential parts of parallel loops by packing and unpacking e-graphs for extracted expressions while preserving original code structures. Array references, branching, loops, function calls, and member references are supported under a dataflow abstraction. Once e-graphs have the inputs, we run equality saturation with an arbitrary set of rewriting rules and succeedingly select optimal solutions featuring minimal total costs according to our model. With new expressions, we update user code while arranging the order of computation and continue to the compiler invocation, since generated code is compatible with NVHPC, GCC, and Clang.
[…]
4. Program Representation
During compiler optimizations, the program semantics and behavior are preserved while improving its performance. Maintaining the order of data accesses is especially important to ensure reproducibility of results. To express clear dependencies among statements, program analysis often relies on static single-assignment form (SSA), which allows only one definition per variable (Appel, 1997). Early work on equality saturation (Tate et al., 2009) used an SSA-based graph structure to represent and rewrite an entire program. For directive/based code, however, user-specified parallelism significantly affects performance (Montgomery et al., 2015), so compilers are limited to respect users’ decisions.
Our proposed methodology optimizes OpenACC/ OpenMP code while preserving a connection between e-graphs and original code. We track optimal expressions in e-graphs using SSA variables and insert optimized code into programs with the same structures and directives as input. The rest of this section covers the conversion process from directive/based code to e-graphs and explains the solution finding process based on SSA. By doing so, we are able to improve the performance of parallel code without changing its semantics or behavior.
4.1. E-Graph Creation
Both OpenACC and OpenMP contain sequential parts within the innermost parallel loops that may be executed independently across multiple threads and thus optimized for more efficient execution or reduced computation. To attain this, we create an e-graph for each innermost parallel loop as follows: First, we introduce conditional nodes (Havlak, 1994) to represent control structures such as if and for, while merging data flows. Second, we assign an ID to each variable/array assignment or . Third, we update each variable/array load to refer to the latest ID along its data flow. Last, for each assignment or , we assign both the ID and the expression to the same e-class.
[…]
4.2. Code Selection
To select optimal codes in the e-graph, we aim to find an equal expression that corresponds to a set of all assignments. We extract the lowest-cost expression that contains all the e-classes of assignments based on a sequence of IDs. The total cost is calculated as the sum of the cost of each e-class, with common e-classes being counted only once. To attain this, we use linear programming techniques (Forrest and Lougee-Heimer, 2005).
Each assignment is updated based on its ID and the extracted expression. To enable reuse across assignments, the values of common expressions are stored in temporary variables. ACC Saturator allows for customizing the generation order of these temporary variables (Section 6).
5. Optimization with Saturation
In this section, we present our approach to rewriting OpenACC/OpenMP code using equality saturation. Our strategy involves a carefully crafted set of rewriting rules and a cost model that aims to minimize both computation and memory access.
5.1. Rewriting Rules
[…]
We apply two sets of rewriting rules to our equality saturation process. The first set introduces fused multiply/adds (FMA) operations, which can improve code generation and resource utilization on GPUs (Rawat et al., 2018). Table 1 lists the minimum set of rules we use. When we encounter expressions that match the FMA pattern, we add the corresponding operations to the e-classes of matched expressions. The second set of rules benefits from the commutative and associative properties of the plus and multiply operators to reorder computation. This can enable common subexpression elimination and produce new FMA operations. We also incorporate constant folding of arithmetic operations with integer and floating-point numbers.
While ACC Saturator can rewrite subtraction, division, memory access order, conditional expressions, and iterations, these rules can increase the size of e-graphs and lead to slow extraction of optimal solutions in real-time. Therefore, we restrict the tool to only use the set of rules mentioned earlier for efficient performance.
| Name | Pattern | Result | |
| FMA1 | A + B * C | FMA(A, B, C) | |
| FMA2 | A - B * C | FMA(A, -B, C) | |
| FMA3 | B * C - A | FMA(-A, B, C) | |
| COMM-ADD | A + B | B + A | |
| COMM-MUL | A * B | B * A | |
| ASSOC-ADD1 | A + (B + C) | (A + B) + C | |
| ASSOC-ADD2 | (A + B) + C | A + (B + C) | |
| ASSOC-MUL1 | A * (B * C) | (A * B) * C | |
| ASSOC-MUL2 | (A * B) * C | A * (B * C) |
5.2. Cost Model
GPUs are complex systems that require careful analysis to accurately predict their efficiency (Khairy et al., 2020; Wu et al., 2015). While applications running on GPUs often face memory-bound limitations (Vijaykumar et al., 2015), it is not just the number of memory accesses that affects overall bandwidth. Factors such as on-chip resource utilization, processor occupancy, thread/ grid-level parallelism across multiple memory layers, and instruction-level parallelism (ILP) all play a role, and improving one metric often comes at the expense of another. In ACC Saturator, our focus is on reducing memory access and computation by utilizing registers for common expressions, while maintaining ILP as described in Section 6.
Our cost model is simple: constant numbers pose no cost, each input variable or counts as 1, all computational operations except division and modular arithmetic count as 10, and each memory access, division, modular arithmetic, or function call counts as 100. The assigned costs are based on empirical testing and aim to reflect the relative cost of different operations. We acknowledge that future research could refine the cost values.
6. Code Generation
To implement the extracted solutions from e-graphs in the input code, we introduce temporary variables. During the generation step, ACC Saturator leverages a novel technique called bulk load that prioritizes high memory pressure by reordering computations.
6.1. Temporary-Variable Insertion
[…]
The compilers of directive-based code can optimize the redundant use of registers. Our code-generation style reduces duplicate computation and leverages optimal instructions, such as FMA, while preserving ILP.
6.2. Bulk Load
As GPUs suffer high memory-access latency, reducing only memory accesses or computation may not lead to improved performance. ACC Saturator follows a different approach by reordering statements to increase memory pressure first and then minimizing memory operations for the remaining execution.
| Name | Compute | Access | Num. | Original Time | |
| Kernels | NVHPC | GCC | |||
| BT | CFD | Halo (3D) | 46 | 14.85s | 28.04s |
| CG | Eigenvalue | Irregular | 16 | 1.27s | 26.17s |
| EP | Random Num | Parallel | 4 | 2.65s | 3.35s |
| FT | FFT | All-to-All | 12 | 3.06s | 3.10s |
| LU | CFD | Halo (3D) | 59 | 15.36s | 24.86s |
| MG | Poisson Eq | Long & Short | 16 | 0.79s | 0.79s |
| SP | CFD | Halo (3D) | 65 | 10.00s | 12.00s |
We address the high memory-access latency on GPUs by utilizing our proposed technique, bulk load. This technique relocates every memory load to the first place where its dependencies are resolved. When multiple loads share one location, we sort these based on their static indices. We prioritize increasing memory pressure at the beginning of the execution and then avoiding memory operations for the rest of the execution. To illustrate the effectiveness of this technique, we compare the performance of a time/consuming kernel in the OpenACC version of NAS Parallel Benchmarks’ BT (NPB-BT) before and after optimization by ACC Saturator. Listings 2 and 3 show the original and optimized code, respectively. Despite featuring the same directives and code structure, the optimized code performs all the loads before the first assignment (temp1), and each subsequent store refers to local variables, leading to minimum number of operations while utilizing FMA.
| Name | Compute | Access | Num. | Size | Original Time (ACC) | Original Time (OMP) | ||||
| kernels | NVHPC | GCC | NVHPC | GCC | Clang | |||||
| ostencil | Jacobi | Halo (3D) | 1 | Ref | 3.87s | 10.28s | 7.75s | 107.54s | 34.60s | |
| olbm | CFD | Halo (3D) | 3 | Ref | 7.11s | 13.32s | 7.11s | 13.47s | 5.91s | |
| omriq | MRI | Structure-of-arrays | 2 | Ref | 16.02s | 16.18s | 5.99s | 18.54s | 11.87s | |
| ep | Random Num | Parallel | 5 | Ref / Test | (CLASS D / W) | 45.33s | 69.91s | 62.42s | 90.35s | 71.32s |
| cg | Eigenvalue | Irregular | 16 | Ref | (> CLASS C) | 4.28s | 662.58s | 5.06s | 19.03s | 18.42s |
| csp | CFD | Halo (3D) | 68 | Ref / Test | (CLASS C / S) | 7.71s | 27.26s | 111.79s | 589.87s | 105.75s |
| bt | CFD | Halo (3D) | 50 | Ref / Test | (CLASS B / W) | 3.24s | 130.43s | 555.44s | 60.45s | 562.83s |
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
7. Experimental Methodology
We implement ACC Saturator in Racket (Felleisen et al., 2015) using XcodeML (CCS), 2009) to parse and generate OpenACC/OpenMP source codes in C. We use the egg library (Willsey et al., 2021) to perform equality saturation. ACC Saturator integrates with NVHPC (NVIDIA Corporation, 2023) and GCC (GNU Project, 2023) for OpenACC/OpenMP compilation, and with Clang (The LLVM Project, 2023) for OpenMP compilation. For evaluation, we use NVHPC 22.9 with options "-O3 -gpu=fastmath -Msafeptr -(acc|mp)=gpu", GCC 12.2.0 with "-O3 -ffast-math -f(openacc|openmp)" and Clang 15.0.3 with "-O3 -ffast-math -fopenmp". Our experiments run on an NVIDIA A100-PCIE-40GB GPU with an Intel Xeon Silver 4114 CPU.
We evaluate the kernel-execution performance of ACC Saturator on two benchmark suites: NAS Parallel Benchmarks in OpenACC/C (NPB) (Xu et al., 2014) and the SPEC ACCEL benchmark suite in both OpenACC/C and OpenMP/C (SPEC) (Corporation, 2014). Tables 2 and 3 provide the detail of NPB and SPEC, respectively. To ensure adequate memory usage, we select CLASS C as the problem sizes of all NPB benchmarks (the largest size within standard test problems), SPEC uses the referential sizes (Ref) except for GCC’s OpenACC cases of ep, sp, and bt, for which we select the testing sizes (Test) due to high execution latency. NPB’s BT, CG, EP, and SP feature the same computation as SPEC’s bt, cg, ep, and csp, but the implementation of NPB is based on OpenACC’s parallel directive while that of SPEC’s OpenACC benchmarks is on the kernels directive. We report the best kernel performance of three executions. To avoid producing incorrect results, we remove the user-specific parallelism in NPB’s CG for GCC, disable the device-side reduction of GCC’s OpenACC in NPB’s LU and MG, and omit the degree specification of the worker parallelism from GCC’s NPB cases, since it surpasses GCC’s thread limit.
On average, ACC Saturator spends 91.8 ms ( = 253.3; 1.4 1885.0 ms) to construct SSA and generate code for each kernel. Equality saturation is executed within a limit of 10,000 e-nodes, 10 seconds of saturation time, 10 rewriting iterations, and a 30-second extraction time limit. The results of the benchmarks show that each kernel requires an average of 0.63 sec ( = 3.37; 0.00 31.2 sec) for equality saturation.
8. Evaluation
| NVHPC (14.85 sec) | + CSE (14.77 sec) | + CSE+SAT (14.75 sec) | + CSE+BULK (12.59 sec) | + ACCSAT (12.23 sec) | |||||||||||||||||||||
[…]
| 13.6% | 10.08 | 78.0 | 34.50% | 152 | 0.19 | 10.05 | 100% | +0.68% | -2 | +0.00 | 10.04 | 99% | +0.22% | +10 | +0.00 | 6.40 | 106% | +38.00% | +103 | -0.06 | 6.57 | 104% | +37.26% | +103 | -0.06 |
| 13.6% | 10.07 | 78.0 | 34.58% | 152 | 0.19 | 10.06 | 100% | -0.74% | -2 | +0.00 | 10.05 | 99% | +0.12% | +10 | +0.00 | 6.40 | 106% | +38.02% | +103 | -0.06 | 6.56 | 104% | +37.26% | +103 | -0.06 |
| 13.5% | 9.98 | 78.0 | 35.86% | 152 | 0.12 | 9.96 | 100% | -0.09% | -2 | +0.00 | 9.97 | 99% | -0.07% | +10 | +0.00 | 6.69 | 106% | +32.68% | +103 | +0.00 | 6.72 | 104% | +32.33% | +103 | +0.00 |
| 8.8% | 6.48 | 116.3 | 54.06% | 176 | 0.12 | 6.42 | 100% | +0.02% | +0 | +0.00 | 6.43 | 100% | +0.01% | +2 | +0.00 | 6.85 | 108% | -2.02% | +78 | +0.00 | 6.28 | 99% | +2.02% | +79 | +0.00 |
| 8.1% | 6.01 | 116.3 | 57.11% | 176 | 0.12 | 6.05 | 100% | -0.14% | +2 | +0.00 | 6.06 | 100% | -0.40% | +2 | +0.00 | 7.14 | 108% | -6.69% | +78 | +0.00 | 6.39 | 99% | -0.01% | +79 | +0.00 |
| 8.0% | 5.92 | 116.3 | 57.32% | 176 | 0.12 | 5.96 | 100% | -0.25% | +2 | +0.00 | 5.98 | 100% | -0.57% | +2 | +0.00 | 7.08 | 108% | -5.92% | +78 | +0.00 | 6.33 | 99% | +0.11% | +79 | +0.00 |
| 5.1% | 0.023 | 0.1 | 39.62% | 47 | 0.62 | 0.024 | 103% | -1.04% | +0 | +0.00 | 0.024 | 103% | -0.03% | +0 | +0.00 | 0.018 | 99% | +5.53% | +39 | -0.31 | 0.018 | 99% | +6.86% | +39 | -0.31 |
| 5.0% | 3.66 | 35.6 | 52.23% | 96 | 0.25 | 3.63 | 97% | +1.37% | -2 | +0.00 | 3.61 | 99% | +1.28% | -4 | +0.00 | 3.38 | 97% | +2.74% | +0 | +0.00 | 3.38 | 99% | +3.43% | +12 | +0.00 |
| 4.6% | 3.39 | 34.9 | 52.50% | 90 | 0.25 | 3.06 | 96% | -0.70% | +23 | +0.00 | 3.07 | 98% | -0.77% | +27 | +0.00 | 3.03 | 96% | +0.61% | +23 | +0.00 | 3.00 | 98% | +0.21% | +27 | +0.00 |
| 4.0% | 0.018 | 0.1 | 44.22% | 47 | 0.62 | 0.019 | 103% | +0.54% | +0 | +0.00 | 0.018 | 103% | +0.01% | +0 | +0.00 | 0.016 | 99% | +2.83% | +39 | -0.31 | 0.016 | 99% | +4.37% | +39 | -0.31 |
| GCC (28.04 sec) | + CSE (25.16 sec) | + CSE+SAT (25.13 sec) | + CSE+BULK (12.79 sec) | + ACCSAT (12.74 sec) | |||||||||||||||||||||
[…]
| 20.1% | 28.01 | 226.2 | 21.66% | 134 | 0.19 | 24.29 | 89% | +2.40% | +0 | +0.00 | 24.26 | 89% | +2.43% | +0 | +0.00 | 5.53 | 50% | +41.55% | +121 | -0.06 | 5.51 | 50% | +41.09% | +121 | -0.06 |
| 20.1% | 28.06 | 226.2 | 19.81% | 130 | 0.19 | 23.80 | 89% | +2.98% | +0 | +0.00 | 23.67 | 89% | +3.00% | +0 | +0.00 | 5.59 | 50% | +43.07% | +125 | -0.06 | 5.58 | 50% | +43.41% | +125 | -0.06 |
| 20.0% | 27.93 | 226.2 | 20.95% | 134 | 0.19 | 24.20 | 89% | +2.44% | +2 | +0.00 | 24.16 | 89% | +2.42% | +0 | +0.00 | 5.46 | 50% | +42.48% | +121 | -0.06 | 5.50 | 50% | +42.16% | +121 | -0.06 |
| 5.4% | 7.54 | 97.5 | 56.38% | 98 | 0.25 | 7.54 | 100% | +0.06% | +0 | +0.00 | 7.55 | 99% | +0.49% | +0 | +0.00 | 6.74 | 101% | +12.70% | +157 | -0.12 | 6.78 | 100% | +12.17% | +157 | -0.12 |
| 5.4% | 7.53 | 97.5 | 56.42% | 98 | 0.25 | 7.53 | 100% | -0.04% | +0 | +0.00 | 7.54 | 99% | -0.02% | +0 | +0.00 | 6.74 | 100% | +12.35% | +157 | -0.12 | 6.77 | 100% | +12.29% | +157 | -0.12 |
| 5.4% | 7.54 | 97.5 | 56.77% | 98 | 0.25 | 7.54 | 100% | -0.76% | +0 | +0.00 | 7.54 | 99% | -0.47% | +0 | +0.00 | 6.76 | 100% | +11.93% | +157 | -0.12 | 6.82 | 100% | +12.10% | +157 | -0.12 |
| 3.9% | 0.034 | 0.3 | 43.50% | 56 | 0.56 | 0.034 | 100% | -0.35% | +0 | +0.00 | 0.034 | 100% | -0.44% | +0 | +0.00 | 0.025 | 92% | -9.16% | +48 | -0.31 | 0.025 | 92% | -8.73% | +48 | -0.31 |
| 3.4% | 4.71 | 37.8 | 48.48% | 96 | 0.31 | 3.58 | 87% | -1.79% | +0 | +0.00 | 3.60 | 86% | -3.22% | +0 | +0.00 | 3.58 | 87% | -2.30% | +0 | +0.00 | 3.43 | 85% | +0.24% | +26 | -0.06 |
| 3.3% | 4.64 | 39.9 | 41.44% | 96 | 0.31 | 4.03 | 87% | +8.93% | +8 | -0.06 | 4.03 | 87% | +8.43% | +10 | -0.06 | 3.88 | 87% | +8.11% | +12 | -0.06 | 3.88 | 87% | +7.09% | +10 | -0.06 |
| 3.2% | 4.50 | 39.1 | 49.96% | 96 | 0.31 | 3.56 | 88% | -3.55% | +0 | +0.00 | 3.54 | 86% | -4.09% | +0 | +0.00 | 3.58 | 89% | -4.06% | +0 | +0.00 | 3.43 | 87% | -1.04% | +14 | -0.06 |
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
{tikzpicture}[remember picture,overlay,>=stealth]
[…]
Figure 2 presents the speedup results of NPB using four generated code versions. CSE is a version that eliminates redundant loads without performing equality saturation or bulk load. CSE+SAT and CSE+BULK are CSE with equality saturation and bulk load, respectively. ACCSAT is a default generated code of ACC Saturator, which includes both equality saturation and bulk load. With ACCSAT, NVHPC attains up to 1.21x improvement, while GCC attains up to 2.20x speedup. The CSE version maintains the performance of both compilers, varying the execution efficiency by 0.98x with NVHPC and by 1.03x with GCC on average. CSE+SAT provides NVHPC with an average speedup of 0.86% and improves GCC’s performance by only 0.01%. However, CSE+BULK significantly accelerates memory-intensive applications such as BT, LU, and SP, while most other benchmarks maintain their performance. ACCSAT does not degrade the original performance and attains 2.00% and 0.66% better throughputs than CSE+BULK on NVHPC and GCC, respectively. In total, ACCSAT attains average speedups of 1.10x on NVHPC and 1.29x on GCC.
Table 4 breaks down the top 10 kernels in NPB’s BT, and Figure 3 shows the speedup of each kernel in each version. The speedups of CSE+BULK and ACCSAT are similar to those of CSE and CSE+SAT, respectively. ACCSAT attains up to 2.23x and 5.08x improvements on NVHPC and GCC, respectively, resulting in up to 2.08x and 3.19x memory bandwidth. The top three kernels on NVHPC suffer performance degradation from CSE+BULK to ACCSAT, because ACCSAT spills more registers to global memory, facilitating the reuse of computation. The next three kernels execute 8.3% fewer instructions and around a 1.11x speedup compared to CSE+BULK, since our optimized code clarifies dependencies and reduces both computation and stores. We expect a total speedup of 1.22x when equality saturation is disabled when registers spill.
On NVHPC, CSE increases the latency of CG, resulting in less SM occupancy due to increased register use. However, CSE+SAT and ACCSAT alleviate this register pressure by utilizing FMA. Equality saturation enables EP to become faster by executing 10.53% more FMA operations and 1.67% fewer total operations than CSE. Although other parts of MG suffer from register pressure, one part of the benchmark obtains a 1.14x speedup from CSE+BULK to ACCSAT by increasing the L1 cache hit ratio. Furthermore, FT, LU, and SP benchmarks attain improved memory throughputs and refined total performance with bulk load.
Figure 4 depicts the speedups of SPEC benchmarks for both OpenACC and OpenMP versions. Benchmarks with names starting with "p-" indicate the OpenMP versions. On average, with OpenACC, CSE, CSE+SAT, CSE+BULK, and ACCSAT attain 1.06x, 1.04x, 1.08x, and 1.08x speedups, respectively on NVHPC, while with OpenMP, these attain 1.07x, 1.06x, 1.37x, and 1.47x speedups, respectively. OpenACC’s ostencil sees a 16.7% reduction in SM occupancy due to equality saturation, leading to decreased performance. CSE reduces memory loads by around 50% for olbm and polbm, resulting in improved performance. Bulk load significantly boosts the performance of memory/intensive benchmarks such as csp, bt, pcsp, and pbt, just as in NPB. ACCSAT attains 3.62x speedup in one part of pbt, which executes only one thread-block over nested loops, by eliminating 77.2% memory loads and 50.9% stores. Our optimization decreases operations and reorders memory accesses, being effective for both parallel and sequential iterations.
GCC attains average speedups of 1.16x, 1.16x, 1.48x, and 1.48x for OpenACC using CSE, CSE+SAT, CSE+BULK, and ACCSAT, respectively, and 1.00x, 1.00x, 1.06x, and 1.06x for OpenMP. The original versions of OpenMP result in high register pressure, which limits the benefits of bulk load. Conversely, the initial versions of OpenACC use fewer registers while setting inadequate parallelism, likely due to the immature support of OpenACC’s kernels directive. CSE reduces memory loads by 54.8% in olbm, yielding a 1.32x speedup. For ep, CSE minimizes operations, and bulk load enhances overall memory throughput. Bulk load also benefits cg, csp, bt, pscp, and pbt by mitigating global-memory latency. However, pomriq’s SM occupancy decreases with bulk load and equality saturation, leading to reduced efficiency.
Clang attains average speedups of 1.06x, 1.06x, 1.69x, and 1.66x for OpenMP using CSE, CSE+SAT, CSE+BULK, and ACCSAT, respectively. CSE+SAT refines the performance of postencil, polbm, and pep compared to CSE. ACCSAT further improves pbt over CSE+BULK, while reducing pomriq’s ILP due to decreased register usage. CSE+BULK attains a maximum speedup of 4.84x through optimized memory accesses.
For comparison, Figure 5 provides the speedups of NPB on an NVIDIA A100-SXM4-80GB, which features 1.31x higher memory bandwidth than the A100-PCIE-40GB. The original performance is improved by 5.79% with NVHPC and 4.65% with GCC on average, while most benchmarks preserve the performance changes on the other GPU. We confirm that BT gains further 1.25x and 2.31x speedups by ACCSAT on NVHPC and GCC, respectively. Using the same GPU, CSE+BULK improves the execution by 1.21x and 2.30x, and ACCSAT improves both computation and memory throughputs in BT more for faster memory systems as the latency of computation becomes distinct. With ACCSAT, NVHPC and GCC attain average speedups of 1.11x and 1.31x, respectively. Our technique mitigates the memory latency that results from GCC’s inadequate thread utilization for the kernels directive, resulting in superior performance gains compared to NVHPC.
Figure 6 shows the speedups of SPEC benchmarks on NVIDIA A100-SXM4-80GB. With OpenACC, NVHPC increases the original performance by 7.42% and GCC increases by 3.11% using the GPU on average, while the original performance with OpenMP is increased by 3.33% on NVHPC, -13.3% on GCC, and 1.04% on Clang. Especially, pcg on GCC suffers from the latency of memory barriers, which decreases the original performance by 59.3% and degenerates the execution of optimized codes. CSE+SAT causes a lower L1 cache hit ratio for pep on NVHPC. Overall, ACCSAT obtains average speedups of 1.09x on NVHPC and 1.47x on GCC with OpenACC and average speedups of 1.45x on NVHPC, 1.02x on GCC, and 1.66x on Clang with OpenMP.
9. Related Work
Maximizing architectural utilization is a significant challenge in HPC, requiring application code optimization. Compilers play a crucial role for enhancing performance by applying various code-generation techniques, enabling both generic and architecture-specific optimizations. Numerous programming models and state-of-the-art techniques, especially for GPUs, have been introduced to adapt codes to parallel architectures (Ragan-Kelley et al., 2013; NVIDIA Corporation, 2022; Stulajter et al., 2022; Paszke et al., 2019; Edwards and Sunderland, 2012). Directive-based programming models (Organization, 2011; ARB, 1997) extend sequential languages, allowing complex scientific applications to offload loop iterations to accelerators while maintaining their structures. However, such compilers often rely on the basis of sequential code generation, thus limiting optimization opportunities to general computation (Matsumura et al., 2021; Barua et al., 2018).
Several projects have explored domain-specific or architecture/specific approaches for directive-based code. The CLAW DSL (Clement et al., 2018) provides directives for grid-based algorithms, enabling target-specific optimizations such as spatial blocking while supporting OpenACC/OpenMP code generation. JACC (Matsumura et al., 2021) is an OpenACC runtime framework offering just-in-time kernel compilation with dynamic constant expansion. OptACC (Montgomery et al., 2015) performs runtime parameter searches to optimize OpenACC parallelism. CCAMP (Lambert et al., 2020) interchanges OpenACC and OpenMP, optimizing parallelization for each combination of models and architecture. Barua et al. (Barua et al., 2018) develop an automated OpenACC-kernel optimizer for maximizing ILP through loop unfolding. SAFARA (Tian et al., 2016) fully utilizes register resources to facilitate array reference reuse in OpenACC kernels.
Our ACC Saturator differs in three aspects: (1) automation of optimization through rewriting rules and a cost model-based optimal code selection, (2) integration of bulk load optimization technique for significant GPU memory throughput improvement, and (3) preservation of original code structures while being applicable to both OpenACC and OpenMP without requiring domain-specific information.
Since the introduction of the equality-saturation library egg (Willsey et al., 2021), numerous studies have leveraged it, particularly in the context of GPU computing, for accelerating deep learning applications (Yang et al., 2021; Fu et al., 2023; Shaikhha et al., 2023; Wang et al., 2020; Smith et al., 2021). These works employ rewriting rules for arithmetic expressions, abstract operations, or tensor graphs to optimize convolutions, sparse tensor algebra, or whole tensor operations. Diospyros (VanHattum et al., 2021) synthesizes efficient DSP operations from C code using equality saturation, while Gowda et al. (Gowda et al., 2022) implement a symbolic algebra system with egg for automatic parallelism assignment.
Although the initial work of equality saturation was demonstrated as a Java bytecode optimizer (Tate et al., 2009), recent works using egg focus on program synthesis and code optimization without control statements. Our work is the first to bridge the gap between user code and equality saturation by automatically extracting computation and constructing SSA information for data dependencies. This approach enables novel equality/saturation optimizations for directive-based programming, as detailed in this paper, without requiring further programmer intervention.
Several works propose innovative code optimization approaches. Ben-Nun et al. (Ben-Nun et al., 2019; Ben-Nun et al., 2023) develop data-centric flow graphs to focus on data-oriented optimizations, such as dead memory elimination. Their representation supports collective operations and macroscopic parallelization, enabling optimizations irrespective of calculation. The MLIR framework (Chelini et al., 2021) utilizes multi-level IRs for cooperative domain-specific optimizations. Ginsbach et al. (Ginsbach et al., 2018) define code patterns for performance opportunities, replacing matched code with library or DSL implementations. Additionally, machine learning is gaining popularity for automatic compiler-optimization tuning (Park et al., 2022; Trofin et al., 2021; Haj-Ali et al., 2020). Our tool employs e-graph operations to identify optimal solutions without source code analysis or abstract syntax tree transformations.
Numerous optimization techniques have been developed for GPU computing (Hijma et al., 2023). Rawat et al. (Rawat et al., 2018) reorder stencil computation using heuristic algorithms to alleviate register pressure and increase SM occupancy. Software systolic arrays (Chen et al., 2019) assign global data to GPU threads, computing results by propagating and accumulating partial results through warp-level shuffles across overlapped blocks, improving data locality without shared memory usage. Hong et al. (Hong et al., 2018) utilize lightweight kernel emulation to provide auto-tuners with performance bottleneck feedback. In contrast, ACC Saturator adopts a straightforward rewriting mechanism, enabling its application to complex programs. Moreover, our method is complementary to other code optimizations.
10. Conclusion
In recent years, there has been a surge in the development of automation tools to maximize the potential of GPUs for various applications. Directive-based programming, in particular, has become a popular method for converting sequential code into parallel code through annotations. However, these abstract models may hinder code optimization efforts, since actual applications often have complex data dependencies that cannot be easily extended within sequential code structures. As a result, techniques for targeting sequential code to GPUs have been developed to attain higher throughputs while accommodating the unique features of GPU parallel architectures.
This paper presents ACC Saturator, an equality/saturation framework for directive-based programming models that optimizes code using rewriting rules and a cost model. It is the first framework to bridge the gap between user code and equality saturation optimizations while preserving original code structures and data dependencies. We introduce a novel technique, bulk load, which we enable through our framework, to generate code with intentional high memory pressure. We demonstrate the effectiveness of ACC Saturator on various practical benchmarks using multiple compilers for both OpenACC and OpenMP on a state-of-the-art GPU architecture. Our analysis highlights the significance of memory-access order and computational reordering, which ACC Saturator enables, for achieving significant performance improvements and increased memory throughput.
Acknowledgement
We are funded by the EPEEC project from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 801051 and the Ministerio de Ciencia e Innovación—Agencia Estatal de Investigación (PID2019-107255GB-C21/AEI/10.13039/501100011033). We gratefully acknowledge the support of NVIDIA Solutions Lab who provided us the remote access to their GPU environment. We would like to acknowledge the NVIDIA AI Technology Center (NVAITC) Europe for their valuable help. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525.